OpenStudy (anonymous):

An adult alligator may weigh 350 pounds. A newly hatched alligator weighs only 1/9 pound. The adult weighs how many times the hatchling? I got 350 but I'm not entirely sure if that's right.. or if I solved it right.. What I did is I took 350/1 {Divided by} 1/9 Then I flipped 1/9 making it 9/1 So then I did 350/1 x 9/1 and got 3150/9 and Simplified it to 350.

OpenStudy (igreen):

Just Divide.. 350 / 1/9 We can multiply by the reciprocal of 1/9. 350 * 9 = ?

OpenStudy (igreen):

Your mistake was \(\dfrac{3150}{9}\). You didn't do it right.. you had to multiply the denominators together. \(\dfrac{350}{1} \times \dfrac{9}{1} \rightarrow \dfrac{350 \times 9}{1 \times 1}\)

OpenStudy (anonymous):

@kaela_stanley lol at your solution.

OpenStudy (igreen):

\(1 \times 1 \neq 9\)

OpenStudy (igreen):

Understand? @kaela_stanley

OpenStudy (anonymous):

Ah, yes..lol I got it now. Simple mistake.. haha thanks! @iGreen

OpenStudy (igreen):

Yep, :P \(\dfrac{350}{1} \times \dfrac{9}{1} \rightarrow \dfrac{350 \times 9}{1 \times 1} \rightarrow \dfrac{3150}{1} \rightarrow 3150\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @kaela_stanley

OpenStudy (anonymous):

That's all.